OpenStudy (anonymous):

The velocity v of the flow of blood at a distance "r" from the central axis of an artery of radius "R" is: v = k(R^2^ - r^2^) where k is the constant of proportionality. Find the average rate of flow of blood along a radius of the artery. (use 0 and R as the limits of integration)

OpenStudy (anonymous):

\[v=k(R^2-r^2)\]?

OpenStudy (anonymous):

Yes, sorry

OpenStudy (anonymous):

too damn many R's

OpenStudy (anonymous):

are you supposed to integrate wrt \[r\]?

OpenStudy (anonymous):

I think I'm suppose to use the mean value thm....

OpenStudy (anonymous):

well "average value" means integrate and divide by the length of the path

OpenStudy (anonymous):

Im sorry, average value thm

OpenStudy (anonymous):

and in this case they are telling you to integrate from 0 to R so length of path is just R

OpenStudy (anonymous):

in other words you want \[\frac{1}{R}\int_0^R k(R^2-r^2)dr\]

OpenStudy (anonymous):

the R and r are not variables, right?

OpenStudy (anonymous):

how would I apply (b - a) to this?

OpenStudy (anonymous):

i think you are integrating wrt r

OpenStudy (anonymous):

so in this case k, R are constants. b = R, a = 0 and b-a = R

OpenStudy (anonymous):

i mean that is what it says in the problem you wrote. use 0 and R as the limits of inegration

OpenStudy (anonymous):

when you integrate wrt r you get \[\int _0^R kR^2-r^2 dr = \frac{2kR^3}{3}\]

OpenStudy (anonymous):

and when you divide by R you get \[\frac{2kR^2}{3}\]

OpenStudy (anonymous):

here we are assuming that k is a constant, because it says where k is the constant of proportionality. and also R is a constant, the artery has radius R

OpenStudy (anonymous):

what varies is the "flow of blood at distance r" at least that is the way i read the problem. the integral is relatively striaighforward

OpenStudy (anonymous):

\[\int _0^R kR^2dr\] is just \[R\times kR^2=kR^3\] because here R is constant

OpenStudy (anonymous):

\[\int_0^R kr^2dr=\frac{kr^3}{3}|_0^R = \frac{kR^3}{3}\]

OpenStudy (anonymous):

subtract and get \[\int _0^R k(R^2-r^2)dr = \frac{2kR^3}{3}\]

OpenStudy (anonymous):

giving \[\frac{1}{R}\int_0^R k(R^2-r^2)dr = \frac{2kR^2}{3}\]

OpenStudy (anonymous):

phew....... it seems to make sense. It's just going to a little while to sink in.... Thanks

OpenStudy (anonymous):

hope i am right! yw

OpenStudy (anonymous):

how would I apply (b - a) to this?

OpenStudy (anonymous):

phew....... it seems to make sense. It's just going to a little while to sink in.... Thanks

OpenStudy (anonymous):

I think you are